# A priori estimates for parabolic Monge-Ampère type equations

### v1

@@ Abstract.
We prove the existence and regularity of convex solutions to the first initial-boundary value problem of the parabolic Monge-Ampère equation
{$u_{t}&=\det D^{2}u\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\quad\text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $\phi$ is a smooth function, $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$, and $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary. Our approach can also be used to prove similar results for $\gamma$-Gauss curvature flow with any $0<\gamma\leq 1$.
@@
The second author was supported by NSFC 12141103 and the China Scholarship Council during a visit to the Australian National University.
@@ 1. Introduction
In this paper, we study the first initial-boundary value problem of the parabolic Monge-Ampère equation
{$u_{t}&=\det D^{2}u\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\quad\text{ on }\partial_{p}Q_{T},\end{aligned}\right.$ (1.1)
where $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$, and $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary. Assume that $\phi$ is a smooth function in $\overline{Q}_{T}$ with $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$ and $\phi(\cdot,0)$ uniformly convex. We say a function $f:(x,t)\in\mathbb{R}^{n}\times\mathbb{R}^{+}\to f(x,t)\in\mathbb{R}$, is parabolically convex if it is continuous and convex in $x$, and nondecreasing in $t$. If $f$ is further uniformly convex in $x$, we say $f$ uniformly parabolically convex.
We consider solutions in the following Banach space.
@@
$||w||_{\tilde{C}^{k,\alpha}(\overline{Q})}=\sup_{\begin{subarray}{c}|\gamma|+2s\leqq k\\ (x,t)\in Q\end{subarray}}\left|D^{\gamma}D_{t}^{s}w(x,t)\right|+\sup_{\begin{subarray}{c}|y|+2s=k\\ (x,t),(y,\tau)\in Q\end{subarray}}\frac{\left|D^{\gamma}D_{t}^{s}w(x,t)-D^{\gamma}D_{t}^{s}w(y,\tau)\right|}{\left(|x-y|^{2}+|t-\tau|\right)^{\alpha/2}}.$
@@
Assume that equation (1.1) satisfies the compatibility condition of any order (see Remark 3.2) and the above assumptions. Then there exists a unique smooth, parabolically convex solution $u$ to the first initial-boundary value problem (1.1).
The proof of Theorem 1.2 depends on the a priori global $\tilde{C}^{2,\alpha}$ estimates. Specifically, we have the following theorem:
@@
Let $u\in\tilde{C}^{4}(\overline{Q}_{T})$ be a parabolically convex solution of equation (1.1). Assume that $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$ with $C^{4}$ boundary, $\phi\in\tilde{C}^{4}(\overline{Q}_{T})$, $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$, and $\phi(\cdot,0)$ is uniformly convex. Then it holds
$||u||_{\tilde{C}^{2,\alpha}(\overline{Q}_{T})}\leq C,$
for some positive constants $C$ and $\alpha\in(0,1)$ depending only on $n,\Omega,c_{0},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
We can also prove Theorem 1.2 and 1.3 for Gauss curvature flow with boundary
{$u_{t}&=\frac{\det D^{2}u}{(1+|Du|^{2})^{(n+1)/2}}\quad\quad\text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\quad\quad\quad\quad\quad\ \ \text{ on }\partial_{p}Q_{T}.\end{aligned}\right.$ (1.2)
Monge-Ampère equations figure in various geometric problems such as the Minkowski problem and geometric flows. The elliptic Monge-Ampère equations have received considerable study since the 1970s. One can refer to [10, 29] and references therein for more related works. In a celebrated work [2], Caffarelli, Nirenberg, and Spruck resolved the existence and global $C^{2,\alpha}$ regularity of smooth solutions to the Dirichlet problem of Monge-Ampère equations.
Parabolic Monge-Ampère equations occur in Gauss curvature flows, which were first introduced by Firey in [11] to model the shapes of worn stones. In 1976, Krylov proposed three parabolic analogs of Monge-Ampère equations [20], and the regularity problem of those equations has been studied by many researchers. See [4, 6, 14, 16, 18, 24, 27, 30, 31, 32, 33] for regularity results of smooth and viscosity solutions, [15, 34, 35] for the Liouville problem, and [1, 7, 8] for the Gauss curvature flow with flat sides. In a remarkable work [30], Tso obtained the interior $\tilde{C}^{2,\alpha}$ estimate for the equation
[…]
More regularity results of the operator $\mathcal{M}u=-u_{t}\det D^{2}u$ can be found in [16, 24]. There are also many regularity results regarding the parabolic Monge-Ampère operators $\mathcal{M}u=-u_{t}+(\det D^{2}u)^{1/n}$ and $\mathcal{M}u=-u_{t}+\log\det D^{2}u$. The global $\tilde{C}^{2,\alpha}$ regularity of solutions to the first initial-boundary value problems of these two operators were proved respectively in [18] and [31]. One can also find some other results in [5, 17, 26, 34].
However, the global $\tilde{C}^{2,\alpha}$ regularity for equations (1.1) and (1.2) remains open for these two classical parabolic Monge-Ampère equations. Equation (1.2) was studied in [25], where the author where the author proved the existence of a kind of self-similar solutions and obtained its gradient estimate when $\phi\equiv 0$. While the higher global regularity is open. The main obstacle is the lack of concavity of the equations.
It is worth to point out that our approach to obtaining the global $C^{2}$ and $\tilde{C}^{2,\alpha}$ estimates requires a combination of corresponding boundary estimates for equation (1.1) and interior estimates for equation (1.5), the equation of the Legendre transform of $u$, rather than studying only one of the equations to obtain it. Here the interior estimates of equation (1.5) depends on the concavity of this equation.
@@
Step 1:  Establish the $C^{2}$ estimate on boundary
@@
$\lambda(D^{2}u)\geq C\quad\text{on}\quad\partial_{p}Q_{T}.$ (1.4)
@@
Due to the lack of concavity, we cannot use the maximum principle to control the global maximum of $|D^{2}u|$ by $\sup_{\partial_{p}Q_{T}}|D^{2}u|$. Instead, we apply the estimates in Step 1 to the equation after Legendre transform.
@@
$\sup_{Q_{T}^{*}}|D^{2}U|\leq\sup_{\partial_{p}Q_{T}^{*}}|D^{2}U|,$
which combined with (1.3) and (1.4) yields an upper bound of eigenvalues $\lambda(D^{2}U)$ in $Q^{*}_{T}.$ Then, by duality we have $\lambda(D^{2}u)\geq C~{}\text{in}~{}Q_{T}.$ By (1.3), it follows that
$C^{-1}\leq\lambda(D^{2}u)\leq C\quad\text{in}\quad Q_{T}.$ (1.6)
@@
This paper is arranged as follows. In Section 2, we establish the a priori estimates in $\tilde{C}^{2,\alpha}(\Omega\times[\epsilon,T])$. We then obtain the existence and uniqueness of solutions in Section 3, completing the proof of Theorems 1.3 and  1.2. Finally, we prove the existence of a unique smooth solution to equation (1.2) in Section 4.
@@ 2. The a priori estimate
{$u_{t}&=\det D^{2}u\quad\quad\text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\ \ \ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$ (2.1)
where $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$ and $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary. Assume that $\phi$ is a smooth function in $\overline{Q}_{T}$ with $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$ and $\phi(\cdot,0)$ uniformly convex.
In this section, we suppose that $u\in\tilde{C}^{4}(\overline{Q}_{T})$ is a parabolically convex solution to equation (2.1) and aim to prove the $\tilde{C}^{2,\alpha}$ a priori estimate to $u$. For this purpose, we shall obtain the $C^{2}$ estimate for $u$ on the parabolic boundary first.
@@ 2.1. Step 1: $C^{2}$ estimates on the boundary
$u(\cdot,t)$ is strictly convex for any $t\in[0,T]$, and the following inequalities hold,
$||u||_{C^{1}(\overline{Q}_{T})}\leq C,$ (2.2)
@@
$1/C\leq u_{t}\leq C.$ (2.3)
Here $C$ is a universal positive constant which only depends on $n$, $\Omega$, $c_{0}$, $||\phi||_{C^{2}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
@@
By the convexity assumption of $u$, $\max_{\overline{Q}_{T}}u=\max_{\partial_{p}Q_{T}}u=\max_{\partial_{p}Q_{T}}\phi$. Suppose that $u$ reaches its minimum in $\overline{Q}_{T}$ at point $p\in Q_{T}$. Then for sufficiently small $\epsilon>0$, $u_{\epsilon}=u+t\epsilon$ also attains its minimum at $p_{\epsilon}\in Q_{T}$. Then we have $\partial_{t}u_{\epsilon}(p_{\epsilon})\leq 0$ and $D^{2}u_{\epsilon}(p_{\epsilon})\geq 0$. It gives that
$u_{t}(p_{\epsilon})<0~{}~{}\text{ and }~{}~{}D^{2}u(p_{\epsilon})\geq 0,$
which leads to a contradiction. Thus we have $\min_{\overline{Q}_{T}}u=\min_{\partial_{p}Q_{T}}u=\min_{\partial_{p}Q_{T}}\phi$. Hence $||u||_{C^{0}(\overline{Q}_{T})}\leq||\phi||_{C^{0}(\overline{Q}_{T})}\leq C$.
As for the first derivative estimate of $u$ w.r.t. $t$, note that $1/C\leq u_{t}\leq C$ on $\partial\Omega\times[0,T]$ for some constant $C$ due to the boundary condition. Here $C$ depends on the maximum and minimum value of $\phi_{t}$ on $\partial\Omega\times[0,T]$. On $\Omega\times\{0\}$, $u_{t}=\det D^{2}u=\det D^{2}\phi$ is also bounded away from zero and infinity thanks to the uniform convexity and $C^{2}$ boundedness of $\phi$. Letting $h=u_{t}$ and differentiating the equation (2.1) w.r.t. $t$ on both sides, we have
$h_{t}=hu^{ij}h_{ij}\quad\text{ in }Q_{T}.$
We claim that $h$ reaches its positive maximum and minimum values on $\partial_{p}Q_{T}$. And this results in estimate (2.3). The strict convexity follows directly from (2.3) and smoothness of $\phi$, see [3] for a reference. To prove the claim, suppose that $h$ attains its positive maximum in $Q_{T}$, so is $h_{\epsilon}=h-\epsilon t$ for sufficiently small $\epsilon$. Then at that point, $h_{\epsilon}>0$, $D^{2}h_{\epsilon}\leq 0$ and $\partial_{t}h_{\epsilon}\geq 0$. Consequently we have $h_{t}>0$, $h>0$ and $D^{2}h\leq 0$ at that point. This contradicts the equation of $h$. The proof is the same for the minimum of $h$. Thus we have proved (2.3).
To estimate $|Du|$, it suffices to prove the bound for $|D_{\nu}u|$ on $\partial\Omega\times[0,T]$ for the outward unit normal vector $\nu$, since $|Du|$ is bounded by $|D\phi|$ on $\Omega\times\{0\}$ and tangentially on $\partial\Omega\times[0,T]$. For any fixed $t_{0}\in[0,T]$, take a smooth function $w$ such that $w=\phi(t_{0})$ on $\partial\Omega\times\{t_{0}\}$, and $\det D^{2}w\geq C_{2}$. Note that such $w$ does exist, and one example is to assume $w$ with $(-d_{\partial\Omega}+d^{2}_{\partial\Omega})\delta+\phi$ near the boundary for some positive constant $\delta$ large enough. Then by the classical comparison principle for Monge-Ampère equations, we have $u\geq w$ in $\Omega\times\{t_{0}\}$. Thus $D_{\nu}u\leq D_{\nu}w\leq C$, with $C$ depending on the maximum and minimum values of $\phi_{t}$, $D\phi$ on $\overline{Q}_{T}$, $D^{2}\phi(\cdot,0)$ on $\overline{\Omega}$ and $\Omega$. Till now we have proved formula (2.2). ∎
@@
$1/C\leq\lambda(D^{2}u)\leq C\quad\text{ on }\partial_{p}Q_{T},$ (2.4)
where $\lambda(D^{2}u)$ denotes the eigenvalues of $D^{2}u$ and the constant $C$ depends only on $n,\Omega,c_{0},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
@@
for $x\in\Omega$. Here $C$ depends on $||D^{2}\phi(\cdot,0)||_{L^{\infty}(\Omega)}$ and the uniform convexity of $\phi(\cdot,0)$. Let $(x_{0},t_{0})$ be an arbitrary point on $\partial\Omega\times[0,T]$, w.l.o.g. $x_{0}=0$ and $\partial\Omega$ be parameterized by $x_{n}=\rho(x^{\prime})$ on a small neighborhood of 00 with $\rho(0)=0$, $D\rho(0)=0$ and $x_{n}\geq 0$. We claim that
$C^{-1}I_{n-1}\leq D_{x^{\prime}x^{\prime}}u(0,t_{0})\leq CI_{n-1}\quad\text{ for any }t_{0}\in[0,T].$ (2.5)
@@
$(\nabla_{x^{\prime}}+\nabla\rho\partial_{n})(u-\phi)=0.$ (2.6)
@@
We can obtain the latter inequality in (2.5) from this formula. The former inequality of (2.5) follows the same proof as in [2] if we regard $u_{t}$ as a data function at the time $t_{0}$, and the estimation depends on the maximum and minimum values of $u_{t}$.
@@
$\rho(x^{\prime})=\frac{1}{2}\sum_{\alpha,\beta<n}B_{\alpha\beta}x_{\alpha}x_{\beta}+O(|x^{\prime}|^{3})$ (2.7)
@@
$\mathcal{T}=\partial_{\alpha}+\sum_{\beta<n}B_{\alpha\beta}(x_{\beta}\partial_{n}-x_{n}\partial_{\beta}).$ (2.8)

### v2

@@ Abstract.
We prove the existence and regularity of convex solutions to the first initial-boundary value problem for the parabolic Monge-Ampère equation
{$-u_{t}+\det D^{2}u&=\psi(x,t)\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\,\ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $\psi,\phi$ are given functions, $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$, and $\Omega\subset\mathbb{R}^{n}$ is a uniformly convex domain. Our approach can also be used to prove similar results for the $\gamma$-Gauss curvature flow with any $0<\gamma\leq 1$.
@@ 1. Introduction
The Monge-Ampère equation arises in various geometric problems, including the Minkowski problem and geometric flows. The elliptic Monge-Ampère equations have been extensively studied since the 1970s; we refer readers to [7, 25] and references therein for further details.
Over the past decades, the following parabolic Monge-Ampère equation has been widely investigated:
(1.1) {$-u_{t}+\rho(\det D^{2}u)&=\psi(x,t)\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\,\ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$, and $\Omega$ is a convex domain in $\mathbb{R}^{n}$.
To our best knowledge, the literature on this equation can be categorized into three main cases:
  • $(i)$
    $\rho(s)=s^{\frac{1}{n}}$.
  • $(ii)$
    $\rho(s)=-1/s$ and $\psi\equiv 0$;
  • $(iii)$
    $\rho(s)=\log s$.
In case $(i)$, Ivochkina and Ladyzhenskaia [14] obtained the a priori $\tilde{C}^{2,\alpha}$ estimates of the first initial-boundary value problem (see also [20]) if $\Omega$ and $\phi(\cdot,0)$ are uniformly convex, and if either
(1.2) $\inf_{\partial_{p}Q_{T}}\phi_{t}+\inf_{Q_{T}}\psi>\frac{1}{2}ad^{2}$
or
[…]
where $a=\max\{\sup_{Q_{T}}\psi_{t},0\}$ and $d=\text{diam}\,\Omega$. The a priori $C^{1,\alpha}$ and $W^{2,p}$ estimates for a generalized form of this equation were established in [13, 4]. Boundary regularity results were subsequently obtained by Tang [23]. Caffarelli-Schauder type estimates were derived in [29] under the assumption that $\psi$ is Lipschitz continuous.
For case $(ii)$, the equation was used by Tso [26] to investigate the Gauss curvature flow, where he derived $\tilde{C}^{2,\alpha}$ a priori estimates. The Schauder estimates and $W^{2,p}$ estimates were later established in [28] and [12], respectively. The Jörgens-Calabi-Pogorelov type results on entire solutions were obtained in [11, 31]. Moreover, one can refer to [20, 16] for another type of parabolic Monge-Ampère equation $-u_{t}\det D^{2}u=\Psi$.
In case $(iii)$, Tso [27] studied the first initial-boundary value problem and derived the $\tilde{C}^{2,\alpha}$ a priori estimates. Xiong and Bao [30] considered a generalized form where $\rho(s)=\eta(\log s)$ satisfies $\eta^{\prime}>0$ and $n\eta^{\prime\prime}\leq\eta^{\prime}$ to make the parabolic operator
$P(-u_{t},D^{2}u):=-u_{t}+\rho(\det D^{2}u)$
concave with respect to $D^{2}u$. Schnürer and Smoczyk investigated the Neumann boundary problem for a similar equation in [22].
However, the seemingly simplest scenario $\rho(s)=s$ remains less understood. Unlike the three cases discussed above, the parabolic operator $P(-u_{t},D^{2}u)=-u_{t}+\det D^{2}u$ is not concave. The existence of a global smooth solution of (1.1) in this case is a long-standing open problem due to the lack of concavity.
In a celebrated work [1], Caffarelli, Nirenberg, and Spruck established global $C^{2,\alpha}$ estimates for the Dirichlet problem of elliptic Monge-Ampère equations $\det D^{2}u=\psi$, proving the existence of unique smooth convex solutions. This naturally raises the question of whether the first initial-boundary value problem of parabolic Monge-Ampère equation
(1.4) {$-u_{t}+\det D^{2}u&=\psi(x,t)\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\,\ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
has such global regularity under suitable conditions. In this article, we provide an affirmative answer to this question under conditions (1.3) given by Ivochkina and Ladyzhenskaia [14]. Specifically, we assume that $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary, and $\psi,\phi$ are smooth functions satisfying the following conditions:
  • $(\mathcal{P}1)$
    $\phi_{t}+\psi\geq c_{0}>0$ on $\partial\Omega\times[0,T]$;
  • $(\mathcal{P}2)$
    $\phi(x,0)$ is uniformly convex with respect to $x$;
  • $(\mathcal{P}3)$
    $\psi(x,t)$ is concave with respect to $x$.
In this paper, we consider solutions in the following Banach space.
@@
$||w||_{\tilde{C}^{k,\alpha}(\overline{Q})}=\sup_{\begin{subarray}{c}|\gamma|+2s\leqq k\\ (x,t)\in Q\end{subarray}}\left|D_{x}^{\gamma}D_{t}^{s}w(x,t)\right|+\sup_{\begin{subarray}{c}|\gamma|+2s=k\\ (x,t),(y,\tau)\in Q\end{subarray}}\frac{\left|D_{x}^{\gamma}D_{t}^{s}w(x,t)-D_{y}^{\gamma}D_{\tau}^{s}w(y,\tau)\right|}{\left(|x-y|^{2}+|t-\tau|\right)^{\alpha/2}}.$
@@
Let $u\in\tilde{C}^{4}(\overline{Q}_{T})$ be convex in $x$ and solve equation (1.4) under the above assumptions. Then it holds
$||u||_{\tilde{C}^{2,\alpha}(\overline{Q}_{T})}\leq C,$
for some positive constants $C$ and $\alpha\in(0,1)$ depending only on $n,T,\Omega,c_{0},||\psi||_{\tilde{C}^{2}(\overline{Q}_{T})},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
By the a priori estimate, we obtain the existence of solution to equation (1.4).
@@
Assume that equation (1.4) satisfies the compatibility condition of any order (see Remark 3.2) and the above assumptions. Then there exists a unique smooth solution $u$, convex in $x$, to the first initial-boundary value problem (1.4).
Remark 1.4.
It is worth pointing out that in subsection 3.2, we take examples to show that condition $({\mathcal{P}3})$ is necessary for the existence of convex solutions to problem (1.4).
We can also prove Theorems 1.2 and 1.3 for Gauss curvature flow with boundary
(1.5) {$u_{t}&=\frac{\det D^{2}u}{(1+|Du|^{2})^{(n+1)/2}}\quad\quad\text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\text{ on }\partial_{p}Q_{T}.\end{aligned}\right.$
Equation (1.5), corresponding to this setting, was studied in [21], where the author proved the existence of certain self-similar solutions and obtained gradient estimates in the homogeneous case $\phi\equiv 0$. However, higher-order global regularity for this case remains an open problem.
Our main difficulty in proving Theorem 1.2 lies in proving the interior second-order derivative estimates and the boundary $\tilde{C}^{2,\alpha}$ estimates under the lack of concavity. It is worth pointing out that our approach to obtaining such global estimates requires a combination of corresponding boundary estimates for equation (1.4) and interior estimates for equation (1.8), the equation of the Legendre transform of $u$, rather than studying only one of the equations to obtain them.
@@
Step 1:  Establish the $C^{2}$ estimate on the boundary
@@
(1.7) $\lambda(D^{2}u)\geq C\quad\text{on}\quad\partial_{p}Q_{T}.$
@@
Due to the lack of concavity, we cannot use the maximum principle to control the global maximum of $|D^{2}u|$ by $\sup_{\partial_{p}Q_{T}}|D^{2}u|$. Instead, we apply the estimates in Step 1 to the equation after the Legendre transform.
@@
$\sup_{Q_{T}^{*}}||D^{2}U||\leq\sup_{\partial_{p}Q_{T}^{*}}||D^{2}U||,$
which combined with (1.6) and (1.7) yields an upper bound for the eigenvalues $\lambda(D^{2}U)$ in $Q^{*}_{T}.$ Then, by duality we have $\lambda(D^{2}u)\geq C~{}\text{in}~{}Q_{T}.$ By (1.6), it follows that
(1.9) $C^{-1}\leq\lambda(D^{2}u)\leq C\quad\text{in}\quad Q_{T}.$
@@
This paper is arranged as follows. In Section 2, we establish the a priori estimates in $\tilde{C}^{2,\alpha}(\Omega\times[\epsilon,T])$. We then obtain the existence and uniqueness of solutions in Section 3, completing the proof of Theorems 1.2 and 1.3. Finally, we prove the existence of a unique smooth solution to equation (1.5) in Section 4.
@@ 2. The a priori estimate
(2.1) {$-u_{t}+\det D^{2}u&=\psi(x,t)\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\,\ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$, and $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary. Assume that $\psi,\phi$ are smooth functions satisfying assumptions $(\mathcal{P}1)-(\mathcal{P}3)$.
In this section, we suppose that $u\in\tilde{C}^{4}(\overline{Q}_{T})$ is convex in $x$ and a solution to equation (2.1). We aim to prove the $\tilde{C}^{2,\alpha}$ a priori estimate for $u$. For this purpose, we shall obtain the $C^{2}$ estimate for $u$ on the parabolic boundary first.
Lemma 2.1 (Comparison Principle).
Assume that $w,v\in\tilde{C}^{2}(\overline{Q}_{T}\setminus\partial_{p}Q_{T})\cap C(\overline{Q}_{T})$ and one of $w,v$ is convex in $x$. If $-w_{t}+\det D^{2}w\geq-v_{t}+\det D^{2}v$ in $\overline{Q}_{T}\setminus\partial_{p}Q_{T}$ and $w\leq v$ on $\partial_{p}Q_{T}$, then $w\leq v$ in $Q_{T}$.
Proof.
See [20, Theorem 14.1]. ∎
@@ 2.1. Step 1: $C^{2}$ estimates on the boundary
$u(\cdot,t)$ is strictly convex for any $t\in[0,T]$, and the following estimates hold,
(2.2) $||u||_{C^{1}(\overline{Q}_{T})}\leq C,$
@@
(2.3) $1/C\leq u_{t}+\psi\leq C.$
Here $C$ is a universal positive constant that depends only on $n,T$, $\Omega$, $c_{0}$, $||\phi||_{\tilde{C}^{2}(\overline{Q}_{T})}$, $\|\psi\|_{L^{\infty}({Q}_{T})}$, $\|\psi_{t}\|_{L^{n+1}({Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
@@
By the comparison principle, we have
±$\pm u\leq\Big{(}\sup_{\partial_{p}Q_{T}}|\phi|+\sup_{{Q}_{T}}|\psi|\Big{)}e^{t}\quad\text{in}\ Q_{T},$
which yields that $\|u\|_{C^{0}(\overline{Q}_{T})}\leq(\sup_{\partial_{p}Q_{T}}|\phi|+\sup_{{Q}_{T}}|\psi|)e^{T}$.
To obtain (2.3), differentiating the equation (2.1) with respect to $t$, we have
(2.4) $\mathcal{L}u_{t}=\psi_{t}\quad\text{ in }\ Q_{T},$
where the linearized operator $\mathcal{L}$ of (2.1) is
(2.5) $\mathcal{L}=-\partial_{t}+\det D^{2}u\cdot u^{ij}\partial_{ij},$
and $u^{ij}$ is the $(i,j)$-entry of the inverse matrix of $D^{2}u$. Since $\psi(x,t)$ is concave in $x$, then
$\mathcal{L}(u_{t}+\psi)=\det D^{2}u\cdot u^{ij}\partial_{ij}\psi\leq 0.$
It follows form the weak maximum principle that
$\inf_{Q_{T}}(u_{t}+\psi)=\inf_{\partial_{p}Q_{T}}(u_{t}+\psi)=\min\{\inf_{\partial\Omega\times(0,T]}(\phi_{t}+\psi),\inf_{\Omega\times\{0\}}(\det D^{2}\phi)\}.$
That is
(2.6) $\inf_{Q_{T}}(u_{t}+\psi)\geq\min\{c_{0},\inf_{\Omega}(\det D^{2}\phi(\cdot,0))\}>0,$
or,
$\inf_{Q_{T}}u_{t}\geq-\sup_{Q_{T}}|\psi|+\min\{c_{0},\inf_{\Omega}(\det D^{2}\phi(\cdot,0)\}.$
By using the parabolic Alexandrov-Bakelman-Pucci maximum principle ([20, Theorem 7.1]) to equation (2.4), we have
[…]
where
$\Gamma_{u_{t}}=\{(x,t)\in Q_{T}:u_{tt}\geq 0,D^{2}u_{t}\leq 0\},\quad C_{n}=(n(n+1)/\omega_{n})^{\frac{1}{n+1}},$
and $\omega_{n}$ is the surface area of unit sphere in $\mathbb{R}^{n}$. Note that on $\Gamma_{u_{t}}$,
$\begin{split}u_{tt}\det(-D^{2}u_{t})&=(\det D^{2}u)^{-\frac{n-1}{n+1}}u_{tt}\cdot\det\Big{[}(\det D^{2}u)^{\frac{2}{n+1}}(D^{2}u)^{-1}(-D^{2}u_{t})\Big{]}\\ &\leq\left(\frac{u_{tt}+(\det D^{2}u)\text{tr}\Big{[}(D^{2}u)^{-1}(-D^{2}u_{t})\Big{]}}{(n+1)(\det D^{2}u)^{\frac{n-1}{n+1}}}\right)^{n+1}\\ &=\left(\frac{-\mathcal{L}u_{t}}{(n+1)(u_{t}+\psi)^{\frac{n-1}{n+1}}}\right)^{n+1}.\end{split}$
It follows that
[…]
Therefore, we have proved (2.3) according to (2.6) and (2.8). The strict convexity follows directly from (2.3) and the smoothness of $\phi$, see [2] for a reference.
To estimate $|Du|$, it suffices to prove the bound for $|D_{\nu}u|$ on $\partial\Omega\times[0,T]$ for the outward unit normal vector $\nu$, since $|Du|$ is bounded by $|D\phi|$ on $\Omega\times\{0\}$ and tangentially on $\partial\Omega\times[0,T]$. For any fixed $t_{0}\in[0,T]$, take a smooth auxiliary function $w(x)$ such that $w=\phi(t_{0})$ on $\partial\Omega$, and $\det D^{2}w\geq\sup_{Q_{T}}(u_{t}+\psi)$ in $\Omega$. Note that such $w$ does exist, and one example is to assume $w$ with $(-d_{\partial\Omega}+d^{2}_{\partial\Omega})\delta+\phi(t_{0})$ near the boundary for some constant $\delta>0$ large enough. Then by the classical comparison principle for Monge-Ampère equations, we have $u\geq w$ in $\Omega\times\{t_{0}\}$. Thus $D_{\nu}u\leq D_{\nu}w\leq C$. Till now we have proved formula (2.2). ∎
@@
(2.9) $1/C\leq\lambda(D^{2}u)\leq C\quad\text{ on }\partial_{p}Q_{T},$
where $\lambda(D^{2}u)$ denotes the eigenvalues of $D^{2}u$ and the constant $C$ depends only on on $n,T$, $\Omega$, $c_{0}$, $||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$, $\|D\psi\|_{L^{\infty}({Q}_{T})}$, $\|\psi\|_{L^{\infty}({Q}_{T})}$, $\|\psi_{t}\|_{L^{n+1}({Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
@@
for $x\in\Omega$. Here $C$ depends on $||D^{2}\phi(\cdot,0)||_{L^{\infty}(\Omega)}$ and the uniform convexity of $\phi(\cdot,0)$. Let $(x_{0},t_{0})$ be an arbitrary point on $\partial\Omega\times[0,T]$, w.l.o.g. $x_{0}=0$ and $\partial\Omega$ be parameterized by $x_{n}=\rho(x^{\prime})$ on a small neighborhood of 00 with $\rho(0)=0$, $D\rho(0)=0$ and $\Omega\subset\{x_{n}>0\}$. We claim that
(2.10) $C^{-1}I_{n-1}\leq D_{x^{\prime}x^{\prime}}u(0,t_{0})\leq CI_{n-1}\quad\text{ for any }t_{0}\in[0,T].$
@@
(2.11) $(\nabla_{x^{\prime}}+\nabla\rho\partial_{n})(u-\phi)=0.$
@@
We can obtain the latter inequality in (2.10) from this formula, as $|Du|,|D\phi|$ and $|D^{2}\phi|$ are bounded. The former inequality of (2.10) follows the same proof as in [1] if we regard $u_{t}+\psi$ as a data function at the time $t_{0}$, and the estimation depends on the maximum and minimum values of $u_{t}+\psi$, $\Omega$ and $\|\phi\|_{\tilde{C}^{4}{(\overline{Q}_{T})}}$.
@@
(2.12) $\rho(x^{\prime})=\frac{1}{2}\sum_{\alpha,\beta<n}B_{\alpha\beta}x_{\alpha}x_{\beta}+O(|x^{\prime}|^{3})$
@@
(2.13) $\mathcal{T}=\partial_{\alpha}+\sum_{\beta<n}B_{\alpha\beta}(x_{\beta}\partial_{n}-x_{n}\partial_{\beta}).$